Solve for x and y by deriving an expression for y from the second equation, and substituting it back into the first equation.
−5x−7y=9
6x+9y=-9

Answer by checkley79(3341)

−5x−7y=9
6x+9y=-9
9y=-6x-9
y=(-6x-9)/9
-5x-7(-6x-9)/9=-9
-5x+(42x+63)/9=-9
(-5x*9+42x+63)/9=-9
(-45x+42x+63)=-9*9
-3x+63=-81
-3x=-81-63
-3x=-144
x=-144/-3
x=48 ans.
y=(-6*48-9)/9
y=(-288-9)/9
y=-297/9
y=33 ans.
Proof:
5*48-7*33=9
240-231=9
9=9